OpenStudy (anonymous):

6z^3-12z^2+18z Show steps

OpenStudy (solomonzelman):

You want to find the roots?

OpenStudy (campbell_st):

I'll write it in expanded form \[3z \times z^2 - 3z \times 4z + 3z \times 6\]

OpenStudy (solomonzelman):

If so, I would do the following: 1. Factor (the entire polynomial) out of \(\large\color{slate}{6z }\). 2. Factor the remaining part, (the one multiplied times \(\large\color{slate}{6z }\).

OpenStudy (anonymous):

I got z(6z^2-12z+18) but that's not right..

OpenStudy (campbell_st):

well there is also a number that is a common factor to all

OpenStudy (campbell_st):

6, -12 , 18 the common factor is ..?

OpenStudy (anonymous):

So when there's a number that is divisible by all of them you divide it? So you would divide by 6.

OpenStudy (campbell_st):

that's correct... so that is taken outside with z... so its

OpenStudy (anonymous):

6z(z^2-6+3)

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

also, why wouldn't you divide the z and the 6.?

OpenStudy (campbell_st):

not quite... 6z is correct but its 6z x -2z = -12z^2 6z x 3 = 18z so its 6z(z^2 -2z + 3)

OpenStudy (campbell_st):

well if you are asked to factor... you need all the factors..

OpenStudy (anonymous):

how did you get the two?

OpenStudy (campbell_st):

the middle term in the original question is -12z^2 so 6z x -2z = -12z^2

OpenStudy (anonymous):

oh yeah nevermine I messed up. 12/6 is 2. thanks.